### v1

# On elliptic and quasiregularly elliptic manifolds

Abstract.
In his book Metric structures for Riemannian and non-Riemannian spaces, Gromov defined two properties of Riemannian manifolds, ellipticity and quasiregular ellipticity, and suggested that there may be a connection between the two. Since then, groups of researchers working independently have proved strikingly similar results about these two concepts. We obtain new topological obstructions to the two properties: most notably, we show that closed manifolds of both types must have virtually abelian fundamental group. We also give the first examples of open manifolds which are elliptic but not quasireguarly elliptic and vice versa. Whether there is a direct connection between these properties—and, in particular, whether they are equivalent for closed manifolds—remains elusive.
1. Introduction
Language barriers between different mathematical subfields often lead to the same ideas being developed independently in different contexts. In this paper, we draw renewed attention to the parallels between elliptic and quasiregularly elliptic manifolds. Both properties are defined by the existence of certain maps from Euclidean spaces; the two types of maps naturally lend themselves to different tools. Although a link between the two notions was already suggested by Gromov in [13, §2.E and 6.D], similar topological obstructions for ellipticity and quasiregular ellipticity have since been developed in parallel, among others by the two authors. The similarity between both the statements and the proofs in the two cases supports Gromov’s intuition.
We summarize the state of the field and prove several new results. For example, we show that elliptic and quasiregularly elliptic closed manifolds must have abelian fundamental group; this follows from known restrictions on the topology of such manifolds. On the other hand, we give a number of examples showing that, for open manifolds, ellipticity and quasiregular ellipticity do not coincide.
The most obvious open question remains: do they coincide for closed manifolds?
1.1. Definitions and their motivations
Let $M$ be an oriented Riemannian $n$-manifold.
A map $f:\mathbb{R}^{n}\to M$ is said to have positive asymptotic degree if
$\limsup_{R\to\infty}R^{-n}\int_{B_{R}(\mathbb{R}^{n})}f^{*}d\operatorname{vol}_{M}>0.$
Gromov [13, 2.41] suggested calling a manifold $M$ elliptic [^1] This terminology is inspired by, but should not be confused with, the rather distinct notion of elliptic spaces from rational homotopy theory. We sometimes say “Gromov-elliptic” to emphasize this difference. if there is a Lipschitz map $f:\mathbb{R}^{n}\to M$ of positive asymptotic degree. One might as well demand that $f$ be $1$-Lipschitz, in which case the map can be thought of as efficiently wrapping $M$ with an infinite sheet of wrapping paper; to be more concise, we will call such a map a wrapping map.
Recently [6, 5], ellipticity has been shown to be crucial to understanding Lipschitz maps to $M$ from other manifolds. For example, given a nullhomotopic $L$-Lipschitz map $N\to M$, is there a $C(L+1)$-Lipschitz nullhomotopy, for some constant $C=C(M,N)$? This turns out to be true if $M$ is elliptic, simply connected, and formal, a condition from rational homotopy theory, and false if $M$ is simply connected and formal but not elliptic. Spaces satisfying all three conditions are called scalable. More general conjectures of Gromov about how the rational homotopy class of a map is constrained by its Lipschitz constant [13, Ch. 7] are likewise confirmed for scalable spaces and false for spaces which are simply connected and formal but not scalable.
A map $f:\mathbb{R}^{n}\to M$ is said to be quasiregular if $f\in W^{1,n}_{\text{loc}}(\mathbb{R}^{n},M)$ and there is a $K>0$ such that for almost every $x\in M$,
$|Df(x)|^{n}\leq K\det Df(x).$
Quasiregular maps are discrete (in the sense that the preimage of every point is a discrete set) and open due to a theorem of Reshetnyak [33]. They should therefore be thought of as generalized branched coverings with local geometric control. A manifold is said to be quasiregularly (or QR) elliptic, a term introduced by Bonk and Heinonen [7] following Gromov’s template, if it admits a quasiregular map from $\mathbb{R}^{n}$.
The motivation for studying quasiregular maps comes from complex analysis, as the quasiregular condition is a relaxation of holomorphicity. Holomorphic maps in one complex dimension satisfy
$|Df(x)|^{2}=\det Df(x).$
In other words, they are conformal except where they have zero derivative. In higher dimensions, Liouville’s theorem shows that maps satisfying this condition are either constant or restrictions of Möbius transformations. Studying quasiregular maps means understanding the degree to which relaxing this condition increases the universe of maps satisfying it.
Finally, a map $f:N\to M$ between metric spaces is said to have bounded length distortion (BLD) if it is locally bilipschitz along paths, that is, there is a constant $L$ such that for all paths $\gamma:[0,1]\to\mathbb{R}^{m}$,
$L^{-1}\operatorname{length}(\gamma)\leq\operatorname{length}(f(\gamma))\leq L\operatorname{length}(\gamma).$
Following Le Donne and Pankka [20], we say that a Riemannian $n$-manifold $M$ is BLD-elliptic if it admits a BLD map from $\mathbb{R}^{n}$. Since BLD maps $\mathbb{R}^{n}\to M$ are both quasiregular and wrapping maps, every BLD-elliptic manifold is both elliptic and quasiregularly elliptic. On the other hand, Pankka has suggested that every quasiregularly elliptic manifold may be BLD-elliptic. This would explain many of the seeming coincidences observed in this paper.
It follows directly from each definition that if $\tilde{M}\to M$ is a covering map, then $\tilde{M}$ is (Gromov-, quasiregularly, BLD-) elliptic if and only if $M$ is. We use this fact without comment below.
1.2. Obstructions to ellipticity and quasiregular ellipticity
We know a number of results that restrict the topology of a closed elliptic or quasiregularly elliptic manifold. Strikingly, all the results we discuss hold for both types, which suggests a possible close relationship between the two definitions.
One of the earliest known obstructions is the following:
Theorem 1.1 (Varopoulos [36, Theorem X.5.1], Gromov [13, Theorem 2.42]).
If $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold, then the growth rate of $\pi_{1}(M)$ is $O(R^{n})$. In particular, $\pi_{1}(M)$ is virtually nilpotent.
The argument uses an isoperimetric inequality. The elliptic case is easy to see. Suppose that $f:\mathbb{R}^{n}\to M$ is a Lipschitz map; it lifts to a map $\tilde{f}:\mathbb{R}^{n}\to\tilde{M}$. Then for every $R$, the volume of $\tilde{f}(\partial B_{R}(\mathbb{R}^{n}))$ with multiplicity is $O(R^{n-1})$. If sets $S\subset\tilde{M}$ satisfy
$\operatorname{vol}_{n}(S)=o(\operatorname{vol}_{n-1}(\partial S)^{n/(n-1)}),$
then the volume with multiplicity of $\tilde{f}(B_{R}(\mathbb{R}^{n}))$ must be $o(R^{n})$ and $f$ cannot be a wrapping map, so $M$ cannot be elliptic. The same statement about isoperimetric inequalities can be shown for quasiregularly elliptic manifolds. The statement about growth follows from a calculation due to Varopoulos.
Gromov gives several other obstructions to ellipticity in [13, §2.E], most of which can be derived from Theorem 1.1 and the following other result:
Theorem 1.2 (Berdnikov–Guth–Manin [5, Theorem 2.3], Heikkilä–Pankka [15]).
Let $M$ be an elliptic or quasiregularly elliptic closed $n$-manifold. Then there is an injective ring homomorphism $H^{*}(M;\mathbb{R})\to\Lambda^{*}\mathbb{R}^{n}$.
In the elliptic case, this is a theorem of Berdnikov, Guth and Manin, who also show a quantitative refinement [5, Theorem 2.3]. The quasiregular result is a refinement of Prywes’ main theorem in [31]; after we began our work an independent proof was posted by Heikkilä and Pankka [15]. We give relatively concise and similar proofs of both statements by combining ideas from [6] and [31]. In both cases, the main idea is to take pullbacks of arbitrary differential forms generating the cohomology ring and take a scaling limit. While the forms may not obey the same relations as the cohomology classes, their scaling limits do, giving a ring homomorphism from $H^{*}(M;\mathbb{R})$ to an algebra of forms on $\mathbb{R}^{n}$. Evaluating this ring homomorphism at a point gives the desired result.
We give a number of corollaries of this result, the most important of which is the following, answering a question of Gromov [13, Question 2.44]:
Corollary 1.3.
If $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold, then $\pi_{1}(M)$ is virtually abelian.
We show this by studying the cohomology of nilmanifolds. By Theorem 1.1, $\pi_{1}(M)$ is virtually nilpotent, and by Theorem 1.2, $H^{*}(M)$ must contain $H^{*}(\pi_{1}(M)^{\text{ab}})$ as a subalgebra. On the other hand, any noncommutativity kills certain cup products between elements of $H^{1}(M)$. We show this using Nomizu’s theorem, which relates the cohomology of nilpotent groups to Lie algebra cohomology, but one could instead use the Serre spectral sequence.
The analogous result for BLD-elliptic manifolds was obtained using different methods by Le Donne and Pankka [20].
Most known obstructions to ellipticity and quasiregular ellipticity follow from Theorem 1.1, Theorem 1.2, and their corollaries. The main exception is the following:
Theorem 1.4 (Gromov [13, Theorem 2.45], Luisto and Pankka [21]).
Suppose that $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold whose fundamental group is virtually $\mathbb{Z}^{n}$. Then $M$ is aspherical (and therefore finitely covered by a torus).
This theorem is particularly interesting because unlike Theorem 1.2, it gives torsion information. For example, it shows that $T^{n}\mathbin{\#}\Sigma$ is not elliptic or quasiregularly elliptic, where $\Sigma$ is a nontrivial simply connected rational homology sphere. [^2] Such a $\Sigma$ can be constructed as follows. By attaching a $3$-cell to $S^{2}$ along a map of degree $2$, one creates a CW complex $X$ whose only nonzero homology group is $H_{2}(X)\cong\mathbb{Z}/2\mathbb{Z}$. By embedding $X$ in $\mathbb{R}^{7}$ and taking the boundary of a neighborhood of $X$, we obtain a $6$-manifold whose homology is the same as that of $X$ in low dimensions. By Poincaré duality, this manifold is a rational homology sphere. The projection map from $T^{n}\mathbin{\#}\Sigma$ to $T^{n}$ is an isomorphism on rational cohomology.
In contrast, for simply connected spaces, we give the following new result:
Theorem 1.5.
If $M$ is a simply connected (or more generally nilpotent) closed manifold, then whether it is elliptic depends only on its rational homotopy type.
A nilpotent space is a space whose fundamental group is nilpotent and acts nilpotently on its higher homotopy groups. In particular, the higher homotopy groups of a nilpotent space are all finitely generated. On the other hand, for the manifold $T^{n}\mathbin{\#}\Sigma$ discussed above, if $k>1$ is minimal such that $H_{k}(\Sigma)\neq 0$, then
$\pi_{k}(T^{n}\mathbin{\#}\Sigma)\cong\pi_{k}(\widetilde{T^{n}\mathbin{\#}\Sigma})\cong H_{k}(\widetilde{T^{n}\mathbin{\#}\Sigma})\cong\mathbb{Z}^{n}[H_{k}(\Sigma)].$
For example, if there is a map $M\to N$ which induces isomorphisms on all rational homology groups, and $M$ and $N$ are both nilpotent or finitely covered by a nilpotent space, then $M$ is elliptic if and only if $N$ is.
Proof.
Suppose that the $n$-manifold $M$ is elliptic, and let $f:\mathbb{R}^{n}\to M$ be a wrapping map. Suppose that $N$ has the same rational homotopy type as $M$. We use Sullivan’s model of rational homotopy theory (see [35, 11, 9]). The manifolds $M$ and $N$ have the same Sullivan minimal algebra $\mathcal{A}$ and minimal models $m_{M}:\mathcal{A}\to\Omega^{*}M$ and $m_{N}:\mathcal{A}\to\Omega^{*}N$. By the shadowing principle [24, Theorem 4.1], [^3] This is proved in [24] only for simply connected target spaces. The generalization to nilpotent spaces in contained in forthcoming work of Kyle Hansen. the DGA homomorphism $f^{*}m_{M}:\mathcal{A}\to\Omega^{*}\mathbb{R}^{n}$ extends to a DGA homomorphism $\Phi:\mathcal{A}\to\Omega^{*}\mathbb{R}^{n}\times[0,1]$ such that
  1. (i)
    The forms in the image of $\Phi$ have bounded operator norm, and
  2. (ii)
    $\Phi|_{t=1}=g^{*}m_{N}$ for some Lipschitz map $g:\mathbb{R}^{n}\to N$.
In particular, since
$\int_{B(R)}f^{*}d\operatorname{vol}_{M}-\int_{B(R)}g^{*}d\operatorname{vol}_{N}=\int_{\partial B(R)\times[0,1]}\Phi=O(R^{n-1}),$
it follows that $g$ is also a wrapping map. ∎
1.3. Construction of elliptic and quasiregularly elliptic manifolds
In general, wrapping maps are more flexible than quasiregular maps, and therefore it is easier to show that a manifold is Gromov-elliptic than quasiregularly elliptic. For example, this follows directly from the definition:
Proposition 1.6.
Let $M$ and $M^{\prime}$ be elliptic manifolds.
  • The product $M\times M^{\prime}$ is elliptic.
  • If $N$ is a manifold of the same dimension as $M$ and there is a map of positive degree $M\to N$, then $N$ is also elliptic.
A general construction is given in [5] for formal manifolds. A simply connected space is formal in the sense of Sullivan [35, §12] if, informally speaking, its rational homotopy type is the simplest one with its rational cohomology ring. An equivalent informal condition is that there are no nontrivial higher-order rational cohomology operations: the only relations between cohomology classes are given by the cup product. For finite complexes, another useful equivalent condition is as follows: a simply connected finite CW complex $X$ is formal if and only if there is an integer $p>1$ and a map $f_{p}:X\to X$ whose induced map on $H^{k}(X;\mathbb{Q})$, for every $k$, is multiplication by $p^{k}$.
Examples of formal manifolds include all simply connected symmetric spaces, such as spheres, complex projective spaces, and Grassmannians. Products and connected sums of formal manifolds are also formal. The simplest and lowest-dimensional example of a non-formal simply connected manifold is the total space $M$ of a fiber bundle $S^{3}\to M\to S^{2}\times S^{2}$ obtained by pulling back the Hopf fibration $S^{3}\to S^{7}\to S^{4}$ along the degree $1$ map $S^{2}\times S^{2}\to S^{4}$ [9, p. 94].
Theorem 1.7 ([5, Theorem C]).
Let $M$ be a formal simply connected closed $n$-manifold. Then $M$ is elliptic if and only if there is an embedding $H^{*}(M;\mathbb{R})\hookrightarrow\Lambda^{*}\mathbb{R}^{n}$.
In particular, since every simply connected manifold of dimension $\leq 6$ is formal, this gives a complete classification for such manifolds.
We note that if there is an embedding $H^{*}(M;\mathbb{Q})\hookrightarrow\Lambda^{*}\mathbb{Q}^{n}$, then there is a map of positive degree from $T^{n}$ to $M$, and therefore $M$ is trivially elliptic. However, a real embedding as above need not be rationalizable; a 9-dimensional example manifold $M$ is given in [5, Remark 5.2]. A wrapping map of this $M$ must therefore be in some sense aperiodic (but may be almost periodic).
In contrast, we know few examples of quasiregularly elliptic manifolds outside low dimensions. What’s more, all the examples we know factor through a torus; even understanding whether the specific $9$-manifold alluded to above is quasiregularly elliptic would be a breakthrough. In low dimensions, a complete classification is available and coincides with that of elliptic manifolds:
  • In dimension $2$, only $S^{2}$ and $T^{2}$ are quasiregularly elliptic. (In fact, every quasiregular map from $\mathbb{C}$ to a Riemann surface or a domain in $\mathbb{C}$ decomposes into a homeomorphism followed by a holomorphic map, a fact known as Stoïlow factorization.)
  • In dimension $3$, only $S^{3}$, $S^{2}\times S^{1}$, $T^{3}$, and their finite quotients are quasiregularly elliptic due to work of Jormakka [19] and the geometrization theorem.
  • In dimension $4$, Heikkilä and Pankka [15] give a classification up to homeomorphism in the simply connected case: all those manifolds which are not excluded by Theorem 1.2 are homeomorphic to quasiregularly elliptic ones by work of Piergallini and Zuddas [30] because they admit a branched covering by $T^{4}$. Specifically, these are
    #$\#^{r}(S^{2}\times S^{2}),\qquad\#^{s}\mathbb{C}P^{2}\#^{t}\overline{\mathbb{C}P^{2}},\qquad 0\leq r,s,t\leq 3,$
    where, by convention, the connected sum of nothing is $S^{4}$. We extend this classification to all $4$-manifolds in Corollary 3.5: up to finite quotients, the only additional examples are $T^{4}$, $T^{2}\times S^{2}$, and $S^{1}\times S^{3}$.
    Note that this is not a complete classification of which $4$-manifolds are quasiregularly elliptic, since quasiregular ellipticity is only known to be invariant up to PL homeomorphism. Some of the manifolds above are known to admit exotic smooth (and therefore PL) structures [2, 3]; it is not known whether any such exotic $4$-manifolds are quasiregularly elliptic.
In higher dimensions, it is known that $\mathbb{C}P^{n}$ is quasiregularly elliptic, as shown in Prywes’s dissertation [32, Prop. 4.1.3].
1.4. Open manifolds
Thus far we have discussed Gromov- and quasiregular ellipticity for closed manifolds. This restriction is natural because it makes the question topological: for closed manifolds, ellipticity is a homotopy invariant, [^4] This follows from the fact that any homotopy equivalence between compact Riemannian manifolds can be made Lipschitz. and quasiregular ellipticity is invariant at least under PL homeomorphism [15, Proposition B.1]. On the other hand, for open manifolds both properties depend on the metric, and in different ways, as we discuss further in §4 and 5.
In prior works on quasiregular ellipticity such as [28, 29], an open manifold is said to be quasiregularly elliptic if it admits a quasiregular map from $\mathbb{R}^{n}$ with respect to some Riemannian metric. This forces quasiregular ellipticity to be a topological property. Even with this definition, Theorem 1.1 extends to the open manifold case:
Theorem 1.8 (Pankka–Rajala [28, Theorem 1.3]).
If $M$ is a quasiregularly elliptic (closed or open) Riemannian $n$-manifold, then the growth rate of any finitely generated subgroup of $\pi_{1}(M)$ is $O(R^{n})$.
Note that $\pi_{1}(M)$ itself may be infinitely generated (e.g. $\mathbb{Q}$); in that case, this result still forces it to be nilpotent of step at most $\sim\sqrt{n}$.
A drawback of this topological definition is that, for example, it erases the difference between the Euclidean and hyperbolic planes, which are in different conformal classes. In the two-dimensional case, the uniformization theorem allows one to fully classify which Riemannian manifolds are quasiregular, up to isometry.
Moreover, in this sense every open manifold would be elliptic, as it is easy to modify any Riemannian metric on an open manifold to make it elliptic. This gives examples of elliptic open manifolds which cannot be quasiregularly elliptic. The easiest such examples have infinite volume, but we also give a construction of finite-volume examples. To prove a version of Theorem 1.8 for ellipticity, we must impose an additional restriction:
Theorem 1.9.
Suppose that $M$ is a Riemannian $n$-manifold of finite volume, and suppose that its volume form can be written as $\omega+d\alpha$ where $\omega$ is compactly supported and $\alpha$ is bounded. If $M$ is elliptic, then every finitely generated subgroup of $\pi_{1}(M)$ has growth $O(R^{n})$.
The hypothesis holds, for example, for open subsets of closed manifolds, as well as for manifolds of pinched negative curvature and nonpositively curved locally symmetric spaces. The easiest non-examples are infinite strings of surfaces joined by thinner and thinner necks. In future work it will be shown that it is equivalent to $M$ having a nonzero Cheeger constant.
Even then, there are manifolds of dimension $n\geq 3$ which have an elliptic metric satisfying this hypothesis, but no quasiregularly elliptic metric. One such example is $S^{n}\setminus X$ where $X$ is a countable subset of the equatorial $S^{n-2}$. We show in Theorem 5.1 that $S^{n}\setminus S^{n-2}$ (with the round metric) is elliptic, and therefore so is $S^{n}\setminus X$. On the other hand, a generalization of the Rickman–Picard theorem [34] by Holopainen and Rickman [18, Theorem 3.1] shows that $S^{n}\setminus X$ cannot be quasiregularly elliptic in the topological sense above.
Finally, we show in §4 that even for metrics on $\mathbb{R}^{2}$, there is no implication between ellipticity and quasiregular ellipticity. In particular, we give the only known example of a Riemannian manifold which is quasiregularly elliptic, but not elliptic.
1.5. Are closed quasiregularly elliptic manifolds elliptic?
This would be the case if every such manifold were BLD-elliptic, since every BLD map is both quasiregular and Lipschitz. There is not much evidence for this conjecture; Gromov [13, 6.42] sketches a proof, further discussed by Bonk and Heinonen [7, §2], that if there is a quasiregular map $\mathbb{R}^{n}\to M$, then in fact there is a Hölder quasiregular map [^5] The case $M=S^{n}$ was done by Miniowitz [26]., but it’s not clear how one would upgrade this to a Lipschitz quasiregular map.
In general, a quasiregular map need not be even locally Lipschitz: for example, consider the radial stretch map $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined by $f(x)=|x|^{\alpha}\cdot x$, for $\alpha<1$. Moreover, there are examples of quasiregular maps with wild branch sets which are not even topologically conjugate to Lipschitz maps [16]. On the other hand, if one assumes that the branch set (the subset of the domain at which the map is not a local homeomorphism) is contained in a subcomplex of codimension $2$, then the map is always topologically conjugate to a locally Lipschitz one, namely a PL branched cover, due to work of Luisto and Prywes [22]. Thus, even if one wants to show that quasiregularly elliptic manifolds are BLD-elliptic—a more restrictive condition than Gromov ellipticity—the main obstacle is still likely to be the global, rather than local, geometry of the map.
1.6. Are closed elliptic manifolds quasiregularly elliptic?
A priori, this may seem unlikely because quasiregularity is such a restrictive condition (for example, in that quasiregular maps are always open and discrete). However, while it is false for open manifolds as discussed above, the question remains open for closed manifolds.
Gromov suggested in [13, p. 67] that it might be possible to obtain a quasiregular map by somehow optimizing over all Lipschitz maps of positive asymptotic degree: for example, looking at $1$-Lipschitz maps which attain the highest possible degree on every $R$-ball. The example of $S^{2}$ is illustrative. Here, for any choice of rectangle on the equator one gets a quasiregular map which is a cover branched at the preimages of the vertices of the rectangle and sends alternate squares of a checkerboard to the north and south hemisphere. If one gives the squares side length $\pi$, then this map can be made $1$-Lipschitz, and presumably has the highest possible degree on all balls. On the other hand, one can achieve the same degree on balls if one forces two branching points to collide with each other, at which point entire lines of the checkerboard are mapped to the point of collision, and the map is no longer quasiregular. Moreover, at that point one can start moving whole strips of the checkerboard laterally, giving many more degrees of freedom of travel through maps that do not seem to even be close to a quasiregular map. In other words, we have described a large continuous family of optimal maps, of which a non-dense subset is quasiregular.
1.7. Key examples for future work
There are three key examples of spaces whose behavior must be understood to know the full picture.
  1. (1)
    Let $\Sigma$ be an $n$-dimensional simply connected rational homology sphere. For what $12\leq k\leq n-$ 1 is $M=(T^{n-k}\times S^{k})\mathbin{\#}\Sigma$ elliptic or quasiregularly elliptic? This series of examples interpolates between those of Theorem 1.4 and the rational invariance of Theorem 1.5. Note that $M$ cannot be a nilpotent space, as the connected sum gives a copy of the module $H_{p}(\Sigma)[\mathbb{Z}^{n-k}]$ in $\pi_{p}(M)$, where $p$ is the minimal degree such that $H_{p}(\Sigma)\neq 0$.
  2. (2)
    Consider the simplest example of a simply connected non-formal manifold mentioned above. This is the $7$-manifold $M$ which completes the pullback square
[…]
    Is this manifold elliptic? We suspect that it is not, but do not know how to prove this.
    Every elliptic formal $n$-manifold has $L$-Lipschitz self-maps of degree $\geq CL^{n}$, for all sufficiently large $L$. On the other hand, the degree of an $L$-Lipschitz self-map of $M$ is at most $L^{20/3}$ for homological reasons. But the same argument does not give an obstruction to building efficient Lipschitz maps $\mathbb{R}^{7}\to M$; indeed, we have a sketch that such a map can be built with
    $\int_{B(R)}f^{*}d\operatorname{vol}_{M}\geq L^{7}/\log L.$
    Nevertheless, it seems difficult and perhaps impossible to get rid of the logarithmic factor.
  3. (3)
    For all known QR elliptic closed manifolds, the quasiregular map is constructed by factoring through a torus. Therefore, the existence of elliptic manifolds which do not admit positive degree maps from a torus poses a challenge for constructing new quasiregular maps.
    One such space is constructed in [5, Remark 5.2] as follows. Take the real Poincaré duality space
    $Y=(S^{2}\times S^{2})/(x,*)\sim(*,x)\mathbin{\#}2\mathbb{C}P^{2}\mathbin{\#}3\overline{\mathbb{C}P^{2}},$
    where * is a basepoint. The cup product $H^{2}(Y)\times H^{2}(Y)\to H^{4}(Y)$ is the quadratic form $\langle 2,1,1,-1,-1,-1\rangle$, which has discriminant $-2$, and therefore is not rationally equivalent to the quadratic form induced by the cup product $H^{2}(T^{4})\times H^{2}(T^{4})\to H^{4}(T^{4})$. Embed $Y$ in $\mathbb{R}^{10}$. The boundary of a thickening of this embedding is a $9$-manifold with the desired property.
2. Proof of the cohomological obstruction
In this section we prove Theorem 1.2. We start with the elliptic version, first given in [5].
Theorem 2.1.
Let $M$ be a Gromov-elliptic closed $n$-manifold. Then there is an injective ring homomorphism $H^{*}(M;\mathbb{R})\to\Lambda^{*}\mathbb{R}^{n}$.
The basic idea, which carries over into the proof of the quasiregular case, is to choose differential forms generating the cohomology ring of $M$ and pull them back along a Lipschitz map of positive asymptotic degree. As we “zoom out” further and further and rescale, the exact forms corresponding to relations between these forms become smaller and smaller. In the limit, the differential forms factor through a ring homomorphism $H^{*}(M;\mathbb{R})\to\Omega^{*}(\mathbb{R}^{n})$. Restricting to a point where this homomorphism is nontrivial, we get a homomorphism to $\Lambda^{*}\mathbb{R}^{n}$.
For the elliptic case, we use flat differential forms as introduced by Whitney [37, Ch. IX]. These can be thought of as formal limits of smooth forms in the flat norm given by
$\lVert\omega\rVert_{\flat}=\lVert\omega\rVert_{\infty}+\lVert d\omega\rVert_{\infty}.$
Flat forms have a number of relevant properties, for example the pullback of a smooth (or flat) form along a Lipschitz function is a flat form [10, Theorem 3.6]. Following Whitney, we write $\Omega_{\flat}^{*}(M)$ for the complex of flat forms on $M$.
Proof.
Let $w_{1},\ldots,w_{r}$ be a generating set for $H^{*}(M;\mathbb{R})$, where the degree of $w_{i}$ is $a_{i}$, and for each $i$ let $\omega_{i}\in\Omega^{a_{i}}(M)$ be a differential form representing $w_{i}$. Let $p_{1},\ldots,p_{s}\in\mathbb{R}[w_{1},\ldots,w_{r}]$ be a list of polynomials such that as a ring,
$H^{*}(M;\mathbb{R})\cong\mathbb{R}[w_{1},\ldots,w_{r}]/(p_{1},\ldots,p_{s}).$
(Here $\mathbb{R}[w_{1},\ldots,w_{r}]$ denotes the free graded commutative ring in variables of the appropriate degrees.) Then for each $j$, $p_{j}(\omega_{1},\ldots,\omega_{r})$ is an exact form, so we can fix $\alpha_{j}\in\Omega^{b_{j}-1}(M)$ such that
$d\alpha_{j}=p_{j}(\omega_{1},\ldots,\omega_{r}),$
where $b_{j}$ is the (graded) degree of $p_{j}$.
Suppose that $f:\mathbb{R}^{n}\to M$ is a $1$-Lipschitz map of positive asymptotic degree. For every $t>0$ define $f_{t}(x)=f(tx)$; this is a $t$-Lipschitz map. Now we consider forms
$\omega_{i,t}=\frac{f_{t}^{*}\omega_{i}}{t^{a_{i}}},\qquad\alpha_{j,t}=\frac{f_{t}^{*}\alpha_{j}}{t^{b_{j}}}.$
Since pulling back along a $t$-Lipschitz map multiplies the infinity-norm of a $k$-form by at most $t^{k}$, we have
$\lVert\omega_{i,t}\rVert_{\infty}\leq 1,\qquad\lVert\alpha_{j,t}\rVert_{\infty}\leq 1/t.$
By definition of positive asymptotic degree, there is an $\varepsilon>0$ and a sequence of $t\to\infty$ such that $t^{-n}\int_{B_{1}(\mathbb{R}^{n})}f_{t}^{*}d\operatorname{vol}_{M}\geq\varepsilon$. By the Arzelà–Ascoli theorem, this sequence has a subsequence $t_{1},t_{2},\ldots\to\infty$ for which the $\omega_{i,t_{k}}$ converge in the flat norm; we have
$\lim_{k\to\infty}\omega_{i,t_{k}}=\omega_{i,\infty}\in\Omega^{a_{i}}_{\flat}(\mathbb{R}^{n}),\qquad\lim_{k\to\infty}\alpha_{j,t_{k}}=0.$
Since flat limits preserve wedge products [37, §IX.17] and differentials (by definition), this implies that the ring homomorphism $\mathbb{R}[w_{1},\ldots,w_{r}]\to\Omega^{*}_{\flat}(\mathbb{R}^{n})$ defined by $w_{i}\mapsto\omega_{i,\infty}$ passes to a well-defined map on the quotient ring $H^{*}(M,\mathbb{R})\cong\mathbb{R}[w_{1},\ldots,w_{r}]/(p_{1},\ldots,p_{s})$, giving a ring homomorphism
$\varphi_{\infty}:H^{*}(M;\mathbb{R})\to\Omega^{*}_{\flat}(\mathbb{R}^{n}).$
Moreover, flat convergence implies that
$\int_{B_{1}(\mathbb{R}^{n})}\varphi_{\infty}(d\operatorname{vol}_{M})\geq\varepsilon.$
In particular, $\varphi_{\infty}(d\operatorname{vol}_{M})$ is nonzero on some set of positive measure. While flat forms are not well-defined pointwise, they are well-defined up to a measure zero set, so we can choose representatives and then choose a point in this set of positive measure where these representatives actually restrict to a ring homomorphism
$H^{*}(M;\mathbb{R})\to\Lambda^{*}\mathbb{R}^{n}.$
This homomorphism sends the fundamental class to a nonzero element, so by Poincaré duality it is injective. ∎
The proof of the quasiregular case uses a similar overall strategy, but is somewhat more complicated, using ideas from [31]. The main idea is still to compare forms defined on larger and larger balls, so that in the limit primitives go to zero. However, the scaling factor is no longer related to the size of the ball; instead, the existence of a useful scaling factor is guaranteed by quasiregularity. Moreover, the sequence converges in an $L^{p}$ rather than a flat sense.
Before we begin the proof, we set up the main tools we will use. As before, let $w_{1},\ldots,w_{r}$ be a generating set for $H^{*}(M;\mathbb{R})$, where the degree of $w_{i}$ is $a_{i}$, and for each $i$ let $\omega_{i}\in\Omega^{a_{i}}(M)$ be a differential form representing $w_{i}$. Let $p_{1},\ldots,p_{s}\in\mathbb{R}[w_{1},\ldots,w_{r}]$ be a list of polynomials such that as a ring,
$H^{*}(M;\mathbb{R})\cong\mathbb{R}[w_{1},\ldots,w_{r}]/(p_{1},\ldots,p_{s}).$
Then for each $j$, $p_{j}(\omega_{1},\ldots,\omega_{r})$ is an exact form, so we can fix $\alpha_{j}\in\Omega^{b_{j}-1}(M)$ such that
$d\alpha_{j}=p_{j}(\omega_{1},\ldots,\omega_{r}),$
where $b_{j}$ is the (graded) degree of $p_{j}$.
Suppose that $f\colon\mathbb{R}^{n}\to M$ is $K$-quasiregular. For $B\subset\mathbb{R}^{n}$ a ball, let
$A(B)=\int_{B}J_{f}=\int_{B}f^{*}d\operatorname{vol}_{M}$
be the counting function for $f$. We consider the pullbacks by $f$ of the differential forms $\omega_{i}$ and $\alpha_{j}$. The pullbacks have to be rescaled, following the procedure in [31, Section 4]. If $B_{k}\subset\mathbb{R}^{n}$ is a sequence of balls, we define forms on $B_{1}=B(0,1)$ by
$\omega_{i,k}$ $=\frac{1}{A(B_{k})^{1/p_{i}}}T_{k}^{*}f^{*}\omega_{i}$
$\alpha_{j,k}$ $=\frac{1}{A(B_{k})^{1/q_{j}}}T_{k}^{*}f^{*}\alpha_{j},$
where $p_{i}=n/a_{i}$ and $q_{j}=n/b_{j}$ and $T_{k}(x)=x_{k}+r_{k}x$ when $B_{k}=B(x_{k},r_{k})$. We summarize the main results in [31, Section 4] as applied to our setting.
Lemma 2.2.
There exists a sequence of balls $B_{k}\subset\mathbb{R}^{n}$ for which the following holds. If $a_{i},b_{j}\neq 0,n$, then:
  1. (i)
    There exist forms $\omega_{i,\infty}$ and $\alpha_{j,\infty}$ so that
    $\lim_{k\to\infty}\omega_{i,k}=\omega_{i,\infty}\quad\text{and}\quad\lim_{k\to\infty}d\alpha_{j,k}=d\alpha_{j,\infty},$
    where the limits are in the weak topology of $L^{p_{i}}(B_{1})$ and $L^{q_{j}}(B_{1})$, respectively.
  2. (ii)
    The sequence $\alpha_{j,k}$ converges strongly in $L^{q_{i}}(B_{1})$ to $\alpha_{j,\infty}$ and $\alpha_{j,\infty}\in W^{1,q_{i}}(B_{1})$.
  3. (iii)
    There exist forms $\tau_{i,k}$ and $\tau_{i,\infty}$ so that $d\tau_{i,k}=\omega_{i,k}$, $d\tau_{i,\infty}=\omega_{i,\infty}$ and $\tau_{i,k}$ converges strongly in $L^{p_{i}}(B_{1})$ to $\tau_{i,\infty}\in W^{1,p_{i}}(B_{1})$.
  4. (iv)
    For every ball $B_{k}$, we have that $A(B_{k})\leq C(n)A(\frac{1}{2}B_{k})$, and $\lim_{k\to\infty}A(B_{k})=\infty$.
If $a_{i}=n$, then $\omega_{i}=d\operatorname{vol}_{M}$ and $\omega_{i,\infty}=d\operatorname{vol}_{\mathbb{R}^{n}}$ on $B_{1}$.
Lemma 2.3.
Up to taking a subsequence of the index $k$, the sequence $\alpha_{j,k}\to 0$ in $L^{\frac{n}{b_{j}-1}}(B_{1})$ and $d\alpha_{j,k}\to 0$ weakly in $L^{q_{j}}(B_{1})$ when $b_{j}\neq n$ and in distribution when $b_{j}=n$.
Proof.
We first show that $\alpha_{j,k}$ converges to 00:
$\|\alpha_{j,k}\|_{\frac{n}{b_{j}-1}}$ $\leq A(B_{k})^{-\frac{b_{j}}{n}}\left({\int_{B_{k}}|f^{*}\alpha_{j}|^{\frac{n}{b_{j}-1}}}\right)^{\frac{b_{j}-1}{n}}$
$\leq A(B_{k})^{-\frac{b_{j}}{n}}\|\alpha_{j}\|_{\infty}K\left({\int_{B_{k}}J_{f}}\right)^{\frac{b_{j}-1}{n}}$
$=\|\alpha_{j}\|_{\infty}KA(B_{k})^{-\frac{1}{n}},$
which converges to 00 as $k\to\infty$ by Lemma 2.2.
We next show that $d\alpha_{j,k}$ converges in distribution. Suppose that $\phi$ is a test form on $B_{1}$, then
[…]
using integration by parts. This term converges to 00 as $k\to\infty$ by the strong convergence of $\alpha_{j,k}$.
Lastly,
$\|d\alpha_{j,k}\|_{\frac{n}{b_{j}}}$ $\leq K\|d\alpha_{j}\|_{\infty}A(B_{k})^{-b_{j}/n}\left({\int_{B_{k}}J_{f}}\right)^{-b_{j}/n}$
$=K\|d\alpha_{j}\|_{\infty}.$
Therefore the sequence is bounded and when $b_{j}\neq n$ there exists a subsequence that converges weakly. Since the sequence converges to 00 in distribution we arrive at the claim of the lemma. ∎
We next show that wedge products of rescaled forms converge to the wedge product of their limits.
Lemma 2.4.
Suppose $\omega_{k}$ and $\omega_{k}^{\prime}$ are two sequences of $a$-forms and $a^{\prime}$-forms constructed as above, where $a$ and $a^{\prime}$ do not equal 00. Let $p=n/a$ and $p^{\prime}=n/a^{\prime}$. If $\omega_{k}\to\omega_{\infty}$ weakly and $\omega_{k}^{\prime}\to\omega_{\infty}^{\prime}$ weakly in $L^{p}(B_{1})$ and $L^{p^{\prime}}(B_{1})$, repectively, then $\omega_{k}\wedge\omega_{k}^{\prime}\to\omega_{k}\wedge\omega_{k}^{\prime}$ weakly in $L^{n/(a+a^{\prime})}(B_{1})$, if $a+a^{\prime}<n$, and converges in distribution, if $a+a^{\prime}=n$.
Proof.
Let $\phi$ be a smooth test form on $B_{1}$. We know that $\omega_{k}$ is the pullback of a closed form and hence exact on $\mathbb{R}^{n}$. By Lemma 2.2, $\omega_{k}$ and $\omega_{\infty}$ are exact on $B_{1}$. Let $\tau_{k}$ and $\tau_{\infty}$ satisfy $d\tau_{k}=\omega_{k}$ and $d\tau_{\infty}=\omega_{\infty}$. We additionally have that $\tau_{k}\to\tau_{\infty}$ strongly in $L^{p}$. We have
$\left|{\int_{B_{1}}(\omega_{k}\wedge\omega_{k}^{\prime}-\omega_{\infty}\wedge\omega_{\infty}^{\prime})\wedge\phi}\right|$ $=\left|{\int_{B_{1}}d(\tau_{k}\wedge\omega_{k}^{\prime}-\tau_{\infty}\wedge\omega_{\infty}^{\prime})\wedge\phi}\right|$
$=\left|{\int_{B_{1}}(\tau_{k}\wedge\omega_{k}^{\prime}-\tau_{\infty}\wedge\omega_{\infty}^{\prime})\wedge d\phi}\right|,$
by integration by parts. Additionally,
$\left|{\int_{B_{1}}(\tau_{k}\wedge\omega_{k}^{\prime}-\tau_{\infty}\wedge\omega_{\infty}^{\prime})\wedge d\phi}\right|\leq\left|{\int_{B_{1}}(\tau_{k}-\tau_{\infty})\wedge(\omega_{k}^{\prime}\wedge d\phi)}\right|+\left|{\int_{B_{1}}\tau_{\infty}\wedge(\omega^{\prime}_{k}-\omega^{\prime}_{\infty})\wedge d\phi}\right|.$
For the first term, by Hölder’s inequality,
$\left|{\int_{B_{1}}(\tau_{k}-\tau_{\infty})\wedge(\omega_{k}^{\prime}\wedge d\phi)}\right|\leq C(n)\|\tau_{k}-\tau_{\infty}\|_{p}\|\omega_{k}^{\prime}\|_{p^{\prime}}\|d\phi\|_{q}$
where $q^{-1}=1-p^{-1}-(p^{\prime})^{-1}$. This converges to 00 by the strong convergence of $\tau_{k}$ and the boundedness of $\omega_{k}^{\prime}$ in $L^{p^{\prime}}$, where both properties are shown in Lemma 2.2. For the second term, $d\phi\wedge\tau_{\infty}$ is an element of $L^{p^{\prime}/(p^{\prime}-1)}$ since $a+b\leq n$. By the weak convergence of $\omega_{k}^{\prime}$, this term also converges to 00. This gives that $\omega_{k}\wedge\omega_{k}^{\prime}$ converges to $\omega_{\infty}\wedge\omega^{\prime}_{\infty}$ in distribution. If $a+b<n$, by Hölder’s inequality we know that
$\|\omega_{k}\wedge\omega_{k}^{\prime}\|_{\frac{n}{a+a^{\prime}}}\leq\|\omega_{k}\|_{p}\|\omega_{k}^{\prime}\|_{p^{\prime}}.$
Therefore the wedge product converges weakly, up to a subsequence, to some element in $L^{n/(a+b)}(B_{1})$. This element must be $\omega_{\infty}\wedge\omega_{\infty}^{\prime}$ by the convergence result in distribution. Note that even when $a+b=n$, we still have that the limiting form is in $L^{1}(B_{1})$. ∎
Proof of Theorem 1.2.
In order to prove the theorem, we would like to show that the map
$\omega_{i}\mapsto\omega_{i,\infty}$
constructed in Lemma 2.2 factors through a ring homomorphism
$\varphi_{\infty}\colon H^{*}(M,\mathbb{R})\to L^{*}(B_{1}),$
where $L^{*}(B_{1})$ is the ring of differential forms such that a degree $a$-form is in $L^{n/a}(B_{1})$. By Lemma 2.3, $\varphi_{\infty}$ preserves the relations on $H^{*}(M,\mathbb{R})$. By Lemma 2.2, $\varphi_{\infty}$ preserves differentials and by Lemma 2.4, $\varphi_{\infty}$ preserves wedge products. Therefore it is a homomorphism.
We next must show that $\varphi_{\infty}$ is injective. Let $\omega$ and $\omega^{\prime}$ be representatives of Poincaré dual classes on $M$. By Lemma 2.4, $\omega_{k}\wedge\omega_{k}^{\prime}\to\omega_{\infty}\wedge\omega^{\prime}_{\infty}$ in distribution, where the limit is in $L^{1}(B_{1})$. Let $V$ be a positive representative of the fundamental class of $M$. Then $\omega\wedge\omega^{\prime}-V=d\alpha$ and, by Lemma 2.3, $d\alpha_{k}\to 0$ in distribution. It therefore suffices to show that $V_{k}=\frac{1}{A(B_{k})}T_{k}^{*}f^{*}V$ does not converge to 00.
Let $\psi\in C_{c}^{\infty}(B_{1})$ be a positive smooth function that is identically to $1$ on $B_{1/2}$. Let $\psi_{k}=\psi\circ T_{k}$, where $T_{k}(x)=x_{k}+r_{k}x$ and $B_{k}=B(x_{k},r_{k})$. We have that
$\int_{B_{1}}\psi V_{k}$ $=\frac{1}{A(B_{k})}\int_{B_{k}}\psi_{k}J_{f}$
$\geq\frac{1}{A(B_{k})}\int_{\frac{1}{2}B_{k}}J_{f}$
$\geq\frac{1}{C(n)},$
by Lemma 2.2, part (iv). Thus, $V_{k}$ does not converge to 00 in distribution, which proves the claim.
We have shown that $\omega_{i,\infty}\wedge\omega^{\prime}_{i,\infty}$ are equal on $B_{1/2}$ for all choices of $\omega_{i}\in H^{*}(M)$. As in the elliptic setting, there exists a full measure set in $B_{1/2}$ where $\omega_{i,\infty}\neq 0$, for all choices of $\omega_{i}\in H^{*}(M)$. At a point in this full measure set, $\varphi_{\infty}$ restricts to a ring homomorphism. ∎
3. Consequences of the cohomological obstruction
We start by proving Corollary 1.3, restated here:
Corollary 1.3.
If $M$ is a quasiregularly elliptic (or Gromov-elliptic) closed $n$-manifold, then $\pi_{1}(M)$ is virtually abelian.
Proof.
By Theorem 1.1, $\pi_{1}(M)$ is virtually nilpotent. Since covers of (quasiregularly, Gromov/̄)elliptic spaces are elliptic, we can assume by passing to a finite cover that $\pi_{1}(M)$ is nilpotent and torsion-free [8, Lemma 2.42].
Let $\Gamma=\pi_{1}(M)$, and let $G$ be its Mal’cev completion: a simply connected (and therefore contractible) nilpotent Lie group which contains it as a lattice [23]. Then the nilmanifold $N=G/\Gamma$ is a $K(\Gamma,1)$, so there is a map $g:M\to N$ which induces a $\pi_{1}$-isomorphism. The pullback and Theorem 1.2 induce maps
[…]
whose composition is injective on $H^{1}$, and therefore on the subalgebra generated by $H^{1}$. Therefore the subalgebra of $H^{*}(N;\mathbb{R})$ generated by $H^{1}(N;\mathbb{R})$ is isomorphic to $\Lambda^{*}\mathbb{R}^{k}$ where $k=\operatorname{rk}H^{1}(N)$. We will show that this implies that $\Gamma$ is abelian.
Nomizu’s theorem [27] implies that $H^{*}(N;\mathbb{R})$ is isomorphic as a ring to the cohomology of the differential graded algebra (DGA) $\Omega^{*}(G)^{G}$ of $G$-invariant differential forms on $G$. We denote this cohomology by $H^{*}(G)$.
This DGA $\Omega^{*}(G)^{G}$ is also known as the Chevalley–Eilenberg complex of the Lie algebra $\mathfrak{g}$ of $G$. As a ring, $\Omega^{*}(G)^{G}\cong\Lambda^{*}\mathbb{R}^{\dim G}$, since every element is obtained by translating a form on the tangent space of the identity. However, the differential is determined by the group structure of $G$. The tangent space of $G$ at the identity can be identified with $\mathfrak{g}$, and the differential on invariant $1$-forms is dual to its Lie bracket, which can be thought of as a function $\mathfrak{g}\wedge\mathfrak{g}\to\mathfrak{g}$.
Consider the lower central series
$G_{0}=G,\quad G_{1}=[G,G],\quad G_{2}=[G_{1},G],\quad\ldots$
and the corresponding tower of quotients
$\cdots\to Q_{3}\xrightarrow{q_{3}}Q_{2}\xrightarrow{q_{2}}Q_{1}\to 0$
where $Q_{i}=G/G_{i}$. In particular, $Q_{1}\cong H_{1}(N)$ is the abelianization of $G$. On the other hand, (1) extends to a sequence of maps
$H^{*}(Q_{1})\xrightarrow{q_{2}^{*}}H^{*}(Q_{2})\to H^{*}(G)\xrightarrow{\iota g^{*}}\Lambda^{*}\mathbb{R}^{n}$
whose composition must be injective.
We will show that if $G$ is not abelian, that is if $Q_{2}\neq Q_{1}$, then the induced homomorphism $q_{2}^{*}:H^{2}(Q_{1})\to H^{2}(Q_{2})$ cannot be injective, deriving a contradiction. To see this, consider the injective homomorphism
$q_{2}^{*}:\Omega^{*}(Q_{1})^{Q_{1}}\to\Omega^{*}(Q_{2})^{Q_{2}}.$
Since $Q_{1}\cong\mathbb{R}^{r}$ for some $r$, the DGA $\Omega^{*}(Q_{1})^{Q_{1}}$ consists of constant forms and has zero differential. On the other hand, since $q_{2}^{*}$ induces an isomorphism on $H^{1}$, all invariant $1$-forms not in its image must have nontrivial differential. Now notice that $\ker q_{2}$ is the center of $Q_{2}$; on the Lie algebra side, the center is $\ker(q_{2})_{*}\subset\mathfrak{q}_{2}$. In other words, the Lie bracket of vectors in $\mathfrak{q}_{2}$ is determined by their projections to $\mathfrak{q}_{1}$. Since the differential of $1$-forms is dual to this Lie bracket, its image lies in $q_{2}^{*}\Omega^{2}(Q_{1})^{Q_{1}}$. Thus the induced map on $H^{2}$ kills a $(\dim Q_{2}-\dim Q_{1})$-dimensional subspace of $H^{2}(Q_{1})$. ∎
So every elliptic and quasiregularly elliptic manifold has a finite cover whose fundamental group is $\mathbb{Z}^{r}$ for some $r$. In fact, we can use our results to derive further restrictions on the topology of elliptic and quasiregularly elliptic spaces:
Corollary 3.1.
Suppose that $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold and $\pi_{1}(M)$ has rank $k$. Then there is finite cover $\tilde{M}$ of $M$ which has an induced map $\tilde{M}\to T^{k}$ which is surjective on real homology.
Such an $\tilde{M}$ is called an essential manifold.
Proof.
First we construct the induced map. Since $\pi_{1}(M)$ is virtually abelian, we can choose $\tilde{M}$ to be a finite cover such that $\pi_{1}(\tilde{M})\cong\mathbb{Z}^{k}$. Then there is a map $f:\tilde{M}\to T^{k}$ which induces an isomorphism on $\pi_{1}(M)$.
Now consider the pullback of $H^{*}(T^{k})$ along $f$. Since $H^{*}(\tilde{M};\mathbb{R})$ embeds in $\Lambda^{*}\mathbb{R}^{n}$, the image of this pullback is the subalgebra generated by a $k$-dimensional vector subspace of $\Lambda^{1}\mathbb{R}^{n}$. This subalgebra is isomorphic to $\Lambda^{*}\mathbb{R}^{k}$, meaning that $f^{*}$ is injective. Dually, $f$ is surjective on homology. ∎
Example 3.2.
Construct an $n$-manifold $M$, $n\geq 4$, by embedding $(T^{3})^{(2)}$ (the $2$-skeleton of $T^{3}$) in $\mathbb{R}^{n+1}$ and letting $M$ be the boundary of a thickening of this embedding. Then $M$ is not elliptic or quasiregularly elliptic.
Corollary 3.3.
Suppose that $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold. Then $\pi_{1}(M)$ cannot have rank exactly $n-1$.
Proof.
Suppose $\pi_{1}(M)$ has rank $n-1$. By passing to a finite cover, we can assume that $\pi_{1}(M)\cong\mathbb{Z}^{n-1}$. Consider an injective map $i:H^{*}(M;\mathbb{R})\to\Lambda^{*}\mathbb{R}^{n}$. The image of $i$ contains an $(n-1)$-dimensional subspace of $\Lambda^{1}\mathbb{R}^{n}$, without loss of generality $\langle dx_{1},\ldots,d_{n-1}\rangle$. The product $dx_{1}\wedge\cdots\wedge dx_{n-1}$ must pair nontrivially with a $1$-form, which then has to contain a nonzero $dx_{n}$ term. This is a contradiction. ∎
Proposition 3.4.
Suppose that $A\subset\Lambda^{*}\mathbb{R}^{n}$ is a graded subalgebra which satisfies Poincaré duality, and let $x\in A^{1}$. Then $A/(x)$ also satisfies Poincaré duality.
Proof.
Choose a basis $x_{1},\ldots,x_{n}\in\mathbb{R}^{n}$ such that $x=x_{1}$, and write $d\operatorname{vol}=x_{1}\wedge\cdots\wedge x_{n}$.
Let $\bar{a}\in A/(x)$ be a nonzero element of degree $k$. We would like to show that there is a $\bar{b}\in A/(x)$ such that
$\bar{a}\bar{b}=\overline{x_{2}\wedge\cdots\wedge x_{n}}.$
Choose a preimage $a\in A$ of $\bar{a}$, and an element $b_{0}\in A$ such that
$ab_{0}=d\operatorname{vol}.$
Then either $ax_{1}\neq 0$ or $b_{0}x_{1}\neq 0$. (Otherwise both $a$ and $b_{0}$ would be multiples of $x_{1}$ in $\Lambda^{*}\mathbb{R}^{n}$, and their product would be zero.) Suppose first that $ax_{1}\neq 0$. Then there is a $b\in A$ such that $ax_{1}b=d\operatorname{vol}$, and we can choose $\bar{b}$ to be the equivalence class of $b$.
Otherwise, $a=x_{1}a^{\prime}$ for some $a^{\prime}\in\Lambda^{*}\mathbb{R}^{n}$ which is not necessarily in $A$. We show that this cannot happen. Consider the vector space $V$ of elements of $A^{k}$ that have this form. We can choose a vector space $W\subset A^{n-k}$ such that the wedge product defines a nondegenerate bilinear form on $V\otimes W$. In particular, multiplication by $x_{1}$ defines a bijection $W\to x_{1}W$. Now we can similarly build a vector space $V^{\prime}\subset A^{k-1}$ which is Poincaré dual to $x_{1}W$. The map $V^{\prime}\to V$ defined by multiplication by $x_{1}$ is injective, and hence surjective since $V^{\prime}$ and $V$ have the same dimension. Therefore $a$ is contained in the ideal $(x)$ and $\bar{a}=0$, contradicting our assumption. ∎
Corollary 3.5.
Every elliptic or quasiregularly elliptic $4$-manifold with infinite fundamental group is finitely covered by a manifold homeomorphic to $T^{4}$, $T^{2}\times S^{2}$, or $S^{1}\times S^{3}$.
Proof.
Let $M$ be such a manifold. By passing to a finite cover we may assume that $M$ has fundamental group $\mathbb{Z}^{2}$ or $\mathbb{Z}$, or is $T^{4}$ by Theorem 1.4. If it is not $T^{4}$, then from Proposition 3.4 it follows that $M$ has the same real cohomology ring as $T^{2}\times S^{2}$ or $S^{1}\times S^{3}$, and in particular has Euler characteristic zero. By [17, Corollary 6.11.1], $M$ is determined by the total Stiefel–Whitney class of the bundle $S^{2}\to M\to T^{2}$ or $S^{3}\to M\to S^{1}$, respectively. Since this class is of finite order, $M$ is finitely covered by either $T^{2}\times S^{2}$ or $S^{1}\times S^{3}$. ∎
This completes the classification of elliptic and quasiregularly elliptic $4$-manifolds up to finite covers. If $\pi_{1}(M)$ is finite, then $M$ is finitely covered by a simply connected manifold which must come from a finite list of 19 manifolds given in [15, Appendix A].
4. Open surfaces
In this section, we discuss the relationship between ellipticity and quasiregular ellipticity for surfaces. Here quasiregular ellipticity is fully understood due to the uniformization theorem: every Riemannian surface is conformally equivalent to $\mathbb{R}^{2}$, $\mathbb{H}^{2}$, $S^{2}$, or one of their quotients. Those that are quasiregularly elliptic are precisely those that are not conformally hyperbolic.
We will see that this is quite independent of whether the surface, if open, is Gromov-elliptic. Thus we give examples of metrics on the plane that are conformally hyperbolic yet Gromov-elliptic, and that are parabolic (that is, conformally equivalent to the Euclidean plane) and yet not Gromov-elliptic. In the next section, we will see that in fact every open manifold, including every open surface, has a Gromov-elliptic metric.
The main difficulty in analyzing such examples is determining whether they are Gromov-elliptic. To determine whether a surface is quasiregularly elliptic, we use a criterion due to Ahlfors [1]: a metric on the plane is parabolic if and only if $\int_{0}^{\infty}\frac{dr}{L(r)}$ diverges, where $L(r)$ is the circumference of the geodesic circle of radius $r$ centered at some fixed point. A special case of this was highlighted and given a simple proof by Milnor [25]: a rotationally symmetric metric is parabolic if its curvature $K$ satisfies $K\geq-\frac{1}{r^{2}\log r}$ for large radii $r$, and hyperbolic if it satisfies $K\leq-\frac{1+\varepsilon}{r^{2}\log r}$.
4.1. A Gromov-elliptic surface which is conformally hyperbolic
Proposition 4.1.
Let $(\mathbb{R}^{2},g)$ be a Riemannian surface whose metric is symmetric about the origin and has curvature $K|_{x}=-d(x,0)^{-2}$. Then $(\mathbb{R}^{2},g)$ is conformally hyperbolic, but Gromov-elliptic.
Another way of getting a surface with these two properties would be to apply the construction in §5.1 to a punctured torus and take the universal cover.
Proof.
The surface is conformally hyperbolic by Milnor’s result.
Fix a geodesic $\gamma$ starting at the origin, for example the $y$-axis. We first understand the exponential map $\varphi_{R}$ at a point $x=\gamma(R)$, applied to a disk of radius $R/2$ around the origin. The curvature on the image of this map is pinched between $-4R^{-2}$ and $-R^{-2}/4$. Then by Toponogov’s theorem, the exponential map from a Euclidean disk of radius $R/2$ to this disk is $4\sinh(1)$-Lipschitz. On the other hand, the map clearly expands areas everywhere.
We define a wrapping map $f:(\mathbb{R}^{2},g_{\text{Eucl}})\to(\mathbb{R}^{2},g)$ as follows:
  • For $n=1,2,\ldots$, define $f|B((0,2^{2n}),2^{2n-1})$ by
    $f(x,y)=\varphi_{2^{2n-1}}(x,y-2^{2n}).$
  • For all points on the $y$-axis, send $(0,y)\mapsto\gamma(y)$.
  • For all other points, compose with a Lipschitz projection to the boundary of the region on which $f$ is defined thus far.
Clearly the resulting map is Lipschitz and has positive asymptotic degree. ∎
4.2. A parabolic surface which is not Gromov-elliptic
Proposition 4.2.
Let $\Sigma$ be the graph in $\mathbb{R}^{3}$, with the induced metric, of a smooth function $f:\mathbb{R}^{2}\to\mathbb{R}$ defined by
$f(x,y)=\log(\sqrt{a^{2}+b^{2}})\varphi\bigl{(}d\bigl{(}(x,y),\bigl{(}a+{\textstyle\frac{1}{2}},b+{\textstyle\frac{1}{2}}\bigr{)}\bigr{)}\bigr{)}\qquad a,b\in\mathbb{Z},\>(x,y)\in[a,a+1]\times[b,b+1],$
where $\varphi:[0,1]\to[0,1]$ is a smooth, non-increasing bump function supported on $[0,\frac{1}{2})$ and equal to $1$ on $[0,\frac{1}{4}]$. Then $\Sigma$ is parabolic, but not Gromov-elliptic.
Informally, $\Sigma$ is a plane with a lattice of spikes whose height is logarithmic in the distance from the origin. The reason this space is not Gromov-elliptic is roughly that any map of positive asymptotic degree would be forced to have local winding number $\sim\log R$ times around many different spikes; however, these local winding numbers add up to a global “turning number”, forcing the restriction of such a map to a circle of circumference $R$ to make $\gtrsim R^{2}/\log R$ full turns each of which encompasses at least one spike. This is an obvious impossibility.
Proof.
The parabolicity of $\Sigma$ can easily be seen using Ahlfors’ criterion: the circumference of a geodesic circle at radius $R$ is $O(R\log R)$, since it consists of a closed curve of length $O(R)$ in the “flat part” of $\Sigma$ together with loops of length $O(1)$ in $O(R\log R)$ spikes. Since
[…]
the integral of the circumferences diverges.
To see that $\Sigma$ cannot be Gromov-elliptic, consider a $1$-Lipschitz map $f_{R}:B(R)\to\Sigma$ such that $|f_{R}(0)|<R$. Identify points in $\Sigma$ with their projection to the plane. We can assume that $f_{R}$ is smooth and transverse to the curves
$C_{a,b}=\Bigl{\{}(x,y\in[a,a+1]\times[b,b+1]:\varphi(x-a,y-b)=\frac{1}{2}\Bigr{\}}$
for every $a,b\in\mathbb{Z}$.
Let $G$ be the CW complex whose cells are vertices, edges, and squares of the integer grid in $\mathbb{R}^{2}$.
Lemma 4.3.
The function $f_{R}$ extends to a $\pi\sqrt{2}$-Lipschitz function $f:B(R+\log R+2)\to\Sigma$ such that $f|_{\partial B(R+\log R+2)}$ maps to the set
$G^{(1)}=\{x,y\in\mathbb{R}^{2}:x\in\mathbb{Z}\text{ or }y\in\mathbb{Z}\}.$
Proof.
It suffices to build a $\pi\sqrt{2}$-Lipschitz homotopy $h:\partial B(R)\times[0,\log R+2]\to\Sigma$ from $f_{R}|_{\partial B(R)}$ to a map with the desired property. The domain of the homotopy is identified with the annulus in the obvious way.
We build this homotopy in three steps. First, we homotope $f_{R}$ out of the “flat top” $U_{a,b}=B_{1/4}(a+\frac{1}{2},b+\frac{1}{2})$ of each spike. Since $U_{a,b}$ is an open set, its preimage under $f_{R}|_{\partial B(R)}$ is a disjoint union of open arcs. On each such arc $A=(\alpha,\beta)$, we perform a linear homotopy from $f_{R}|_{A}$ to the constant-speed shortest path between the endpoints $f(\alpha)$ and $f(\beta)$ along the boundary of $U_{a,b}$. This homotopy is $\frac{\pi}{2}$-Lipschitz and takes time $1/2$; call its end result $g:\partial B(R)\to\Sigma$.
Next we push the curve down the side of the spike, as with a squeegee: a point starts moving once the homotopy reaches down to its height. This homotopy takes time $<\log(2R)+\frac{1}{4}<\log R+1$ and is $\pi$-Lipschitz, since distances increase by a factor of at most $2$ from $g$.
Finally, we push points from the circle of radius $1/2$ around $(a+\frac{1}{2},b+\frac{1}{2})$ linearly out to the square; at this point we accrue the factor of $\sqrt{2}$, and the homotopy takes time $\sqrt{2}-1$. ∎
In particular, every $f^{-1}(C_{a,b})$ is a closed curve and $f$ has the same local degree over every point in a grid square.
The domain can then be rescaled to make the function once again $1$-Lipschitz. We will show that the resulting function satisfies
$\int_{B(\pi\sqrt{2}(R+\log R+2))}f^{*}d\operatorname{vol}_{\Sigma}=O(R^{2}/\log R),$
and therefore the original function $f_{R}$ likewise has degree $O(R^{2}/\log R)$. Now if $f_{\infty}:\mathbb{R}^{2}\to\Sigma$ is a $1$-Lipschitz map, for large enough $R$ the restriction $f_{R}=f_{\infty}|_{B(R)}$ satisfies the assumption $|f_{R}(0)|<R$. Thus $f_{\infty}$ cannot have positive asymptotic degree.
Now write $B=B(\pi\sqrt{2}(R+\log R+1))$.
Lemma 4.4.
The sum of the local degrees of $f$ over points in
$\Bigl{(}\mathbb{Z}^{2}+\Bigl{(}\frac{1}{2},\frac{1}{2}\Bigr{)}\Bigr{)}\cap[-\sqrt{R},\sqrt{R}]^{2}$
is $O(R^{3/2})$.
Proof.
Consider the homology class induced by $f(\partial B)$ in $G^{(1)}$. This induces a unique cellular cycle $z\in C_{1}(G)$; the mass of $z$ (that is, the sum of the coefficients) is $O(\operatorname{length}(f(\partial B)))=O(R)$. Let $c$ be the (likewise unique) $2$-chain filling $z$; we must estimate the portion of the mass of $c$ contained in $[-\sqrt{R},\sqrt{R}]^{2}$.
Let $c(i)=\{\text{2-cells }\alpha:c_{\alpha}\geq i\}$. Then $\sum_{i=1}^{\infty}\#c(i)\geq\operatorname{mass}(c)$ and
$\operatorname{mass}(z)\geq\sum_{i=1}^{\infty}\#\partial c(i)\geq\sum_{i=1}^{\infty}4\sqrt{\#c(i)},$
where the last inequality holds by the isoperimetric inequality for $(\mathbb{R}^{2},\ell^{\infty})$.
Now let $V$ be the portion of the mass of $c$ contained in $[-\sqrt{R},\sqrt{R}]^{2}$. We have
$V\leq\sum_{i=1}^{\infty}\min\{a(i),4R\}\leq 2\sqrt{R}\sum_{i=1}^{\infty}\sqrt{a(i)}\leq\frac{1}{2}\sqrt{R}\operatorname{mass}(z)=O(R^{3/2}).\qed$
Define the turning number of a non-backtracking path in $G^{(1)}$ as
#(left turns)#(right $\frac{\#(\text{left turns})-\#(\text{right turns})}{4}.$
Every free loop $\gamma:S^{1}\to G^{(1)}$ is homotopic to a unique non-backtracking (“reduced”) loop; we define $t(\gamma)$ to be the turning number of that loop. Then the following is clear:
Lemma 4.5.
  1. (a)
    If $\alpha$ and $\beta$ are two loops, then
    $|t(\gamma)-t(\alpha)-t(\beta)|\leq 1$
    where $\gamma$ is any concatenation of $\alpha$ and $\beta$, that is, any loop that traverses $\alpha$, then travels along a curve $\delta$ from $\alpha$ to $\beta$, then traverses $\beta$, and then travels back along $\delta$. The inequality does not depend on the loop $\delta$.
  2. (b)
    For any loop $\gamma$, $t(\gamma)\leq\operatorname{length}(\gamma)/4$.
Now we slightly modify the concept of turning number to be well-defined for free homotopy classes of curves in $G^{(1)}\cup[-\sqrt{R},\sqrt{R}]^{2}$, which we denote by $t_{\operatorname{sq}}$. Here, we produce a canonical representative by replacing each section of curve that traverses $[-\sqrt{R},\sqrt{R}]^{2}$ by any simple curve between its endpoints. It’s easy to see that this gives rise to a well-defined notion of turning number and that Lemma 4.5 still holds for $t_{\operatorname{sq}}$.
Lemma 4.6.
The sum of the local degrees of $f$ over points in $\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2})$ is $O(R^{2}/\log^{2}R)$.
Proof.
Write
$A=\mathbb{R}^{2}\setminus\Bigl{(}\mathbb{Z}^{2}+\Bigl{(}\frac{1}{2},\frac{1}{2}\Bigr{)}\Bigr{)}\cup[-\sqrt{R},\sqrt{R}]^{2}.$
Since $A$ deformation retracts to $G^{(1)}\cup[-\sqrt{R},\sqrt{R}]^{2}$, we can consider the $t_{\operatorname{sq}}$ numbers of curves in $A$. Let $\gamma_{1},\ldots,\gamma_{s}$ be the components of $f^{-1}C_{a,b}$ for
$\bigl{(}a+{\textstyle\frac{1}{2}},b+{\textstyle\frac{1}{2}}\bigr{)}\notin[-\sqrt{R},\sqrt{R}]^{2}$
which are nontrivial in the sense that they have nonzero winding number around the point $(a+\frac{1}{2},b+\frac{1}{2})$. The turning numbers of these curves are simply these winding numbers. Moreover, $f|_{\partial B}$ is homotopic to a concatenation in $A$ of all the $f\circ\gamma_{i}$, and therefore
$t_{\operatorname{sq}}(f|_{\partial B})\geq\sum_{i}t(\gamma_{i})-\#\{\gamma_{i}\}$
by Lemma 4.5(a). Each nontrivial component winds around a region which contains a preimage $u\in f^{-1}(a+\frac{1}{2},b+\frac{1}{2})$. Since $f$ is $1$-Lipschitz, the region contains a ball of radius $\frac{1}{2}\log R$ around $u$. Therefore, the number of nontrivial components #$\#\{\gamma_{i}\}$ is $O(R^{2}/\log^{2}R)$.
Since $f$ is $1$-Lipschitz, $f_{\partial B}$ has length $O(R)$, so by Lemma 4.5(b), $t_{\operatorname{sq}}(f|_{\partial B})=O(R)$. It follows that $\sum_{i}t(\gamma_{i})=O(R^{2}/\log^{2}R)$. This gives the sum of local degrees of $f$ over points not in $[-\sqrt{R},\sqrt{R}]^{2}$. Combining this with Lemma 4.4, which covers the points inside this square, we get the result. ∎
Grid squares in the image of $f$ have area bounded by $14\log((\pi\sqrt{2}+1)R)+$ 1. Moreover, since we assumed that $f|_{\partial B}$ maps to $G^{(1)}$, $f$ has the same local degree over all points in a grid square. By Lemma 4.6,
$\int_{B}f^{*}d\operatorname{vol}_{\Sigma}\leq(4\log((\pi\sqrt{2}+1)R)+1)O(R^{2}/\log^{2}R)=O(R^{2}/\log R),$
as desired. ∎
4.3. Some remarks on surfaces with nontrivial topology
One application of Theorem 1.9, proved in the next section, is that a Lipschitz map $\mathbb{R}^{2}\to S^{2}$ of positive asymptotic degree can miss at most two points, since $S^{2}$ minus three points has a free fundamental group, which has exponential growth.
Now we show:
Lemma 4.7.
There is a Lipschitz map $f:\mathbb{R}^{2}\to S^{2}$ which has asymptotic degree and misses two points.
This is similar to the case of quasiregular maps—however, as we see in the next section, even in the setting of maps to the sphere missing points, the two cases diverge in higher dimensions.
Proof.
Without loss of generality, the missing points are the north and south pole. As part of the construction, we first give a map
$f_{0}:[-1,1]\times[0,\infty)\to B_{1}(\mathbb{R}^{2})\setminus\{(0,0)\}$
which maps the boundary of the domain to the boundary of the disk. This map is defined explicitly on $[0,1]\times[0,\infty)$ by the polar coordinate formula
$f_{0}(x,y)=(r(x,y),\theta(x,y)),\quad\text{where}\quad\biggl{\{}\begin{aligned} r(x,y)&=x+e^{-y}(1-x)\\ \theta(x,y)&=y+\ln r(x,y).\end{aligned}$
When $x<0$, we define $f_{0}(x,y)=(r(|x|,y),-\theta(|x|,y))$.
$(0,0)$
Figure 1. The domain of the map $f$. The regions outlined in pink and blue map to the northern and southern hemisphere, respectively. The black curves are preimages of the prime meridian, and each region bounded by the black curves is a homeomorphic preimage of the rest of the sphere. The image is not to scale; it is vertically compressed by a factor of $2\pi$.
To define $f:\mathbb{R}^{2}\to S^{2}$, let $p_{+},p_{-}:D^{2}\to S^{2}$ be bilipschitz homeomorphisms of the disk onto the northern and southern hemisphere, respectively, which agree on $\partial D^{2}$. Then we define $f$ everywhere by
$f(x,y)$ $=p_{+}(f_{0}(x,y))$ $(x,y)$ $\in[-1,1]\times[0,\infty)$
$f(x,y)$ $=p_{-}(f_{0}(2-x,y))$ $(x,y)$ $\in[1,3]\times[0,\infty)$
$f(x+4,y)$ $=f(x,y)$
$f(x,-y)$ $=-f(x,y).$
For an illustration of this map, see Figure 1.
A computation shows that:
  • $f$ is Lipschitz;
  • $f$ has the same orientation at every regular point;
  • The Jacobian of $f$ is bounded below by some $\varepsilon>0$ on the set
    $\biggl{(}\bigcup_{z\in\mathbb{Z}}[2z+1/2,2z+3/2]\biggr{)}\times\bigl{(}(-\infty,-1]\cup[1,\infty)\bigr{)}.$
The last two conditions imply that $f$ has positive asymptotic degree. ∎
5. General open manifolds
Now we expand the discussion from surfaces to open manifolds of all dimension. Here we give further constructions of manifolds which are elliptic, but not quasiregularly elliptic. For example, we show that every open manifold has an elliptic metric, and at least a large class (all manifolds with a spherical end) have elliptic metrics of finite volume. A great many of these manifolds admit no quasiregularly elliptic metric. We also prove Theorem 1.9, which proves a restriction on the fundamental group of an open elliptic manifold whose metric satisfies a certain rather strong assumption; this is analogous to a result for quasiregular ellipticity which requires no such assumption. Nevertheless, even open subsets of compact manifolds (for which Theorem 1.9 holds) may be elliptic without being quasiregularly elliptic:
Theorem 5.1.
Let $n\geq 2$, and consider $S^{n-2}$ embedded in the round $S^{n}$ as the intersection with an $(n-1)$-plane through the origin in $\mathbb{R}^{n+1}$. The open manifold $S^{n}\setminus S^{n-2}$ is Gromov-elliptic, and therefore so is $S^{n}\setminus X$ for any $X\subseteq S^{n-2}$.
In contrast, for any countably infinite set $X$, a generalization of the Rickman–Picard theorem [34] by Holopainen and Rickman [18, Theorem 3.1] shows that $S^{n}\setminus X$ does not admit a metric under which it is quasiregularly elliptic.
5.1. An elliptic metric on any open manifold
Let $M$ be any open $n$-manifold and choose any complete Riemannian metric $g$. We will show that the metric $g$ can be modified to construct an elliptic metric, essentially by cutting it open and splicing in a copy of Euclidean space.
Let $\gamma:[0,\infty)\to M$ be a ray, that is, a proper embedding parametrized by arclength. We modify $(M,g)$ by cutting out a small tubular neighborhood of $\gamma([0,\infty))$ and replacing it with an isometric copy of $\mathbb{R}^{n}$ missing a small tubular neighborhood of the negative $x_{1}$-axis. After a bit of smoothing, we get a new Riemannian manifold $(M^{\prime},g^{\prime})$ which is obviously diffeomorphic to $M$ and which includes an isometric copy of $[1,\infty)\times\mathbb{R}^{n-1}$. But then the map
$f(x_{1},\ldots,x_{n})=\begin{cases}(x_{1},\ldots,x_{n})&x_{1}\geq 1\\ (1,x_{2},\ldots,x_{n})&x_{1}\leq 1\end{cases}$
is a wrapping map for $M^{\prime}$.
5.2. A finite-volume elliptic manifold with any fundamental group
Let $(M,g)$ be any Riemannian $n$-manifold of finite volume. We will show that by puncturing $M$ and modifying the metric in a neighborhood of the puncture, we can construct an elliptic manifold whose volume is still finite. In particular, if $\dim M\geq 3$, this preserves the fundamental group. This construction also produces examples of finite-volume elliptic surfaces whose fundamental group is a free group on any number of generators.
Specifically, we build a metric $\tilde{g}$ on $C=S^{n-1}\times[0,\infty)$ by embedding it in $\mathbb{R}^{n+1}$ as a solid of revolution obtained by revolving the graph of a function $\rho(t)$. We define $\rho$ so that for every $p\in\mathbb{N}$, it has a local minimum $\rho(p)=2^{-\frac{2p}{n-1}}$ at $p$, and a local maximum $\rho(p+1/2)=2^{-\frac{p}{n-1}}$; between these, the value of $\rho$ is interpolated smoothly so that the average value over the interval is proportional to the local maximum. In particular, each “nodule” $N_{p}$ from $p$ to $p+1$ has volume proportional to $2^{-p}$, and the total volume of $(C,\tilde{g})$ is finite. In between the nodules, at $p$, the metric has a spherical neck of surface area proportional to $2^{-2p}$.
We now describe a wrapping map $f:\mathbb{R}^{n}\to(C,\tilde{g})$. It will be defined on a sequence of cubes
$Q_{p}=[-2^{p-1},2^{p-1}]^{n-1}\times[2^{p},2^{p+1}].$
These cubes fill a proportion of roughly $2^{-n}$ of any ball around the origin, and thus it is sufficient for the restriction to them to have positive asymptotic degree. On the portion of $\mathbb{R}^{n}$ outside $\bigcup_{p}Q_{p}$, we define $f$ by first projecting to the boundary of $\bigcup_{p}Q_{p}$ (via a $\sqrt{n}$-Lipschitz map) and then applying the restriction to the relevant $Q_{p}$.
Inside $Q_{p}$, we choose a submanifold $S$ isometric to a $(n-1)$-dimensional rectangle which is folded up like an accordion, with $\partial S\subseteq\partial Q_{p}$, and so that the exponential map on its normal $1$-disk bundle is an embedding and fills up a proportion of $Q_{p}$ which depends only on $n$. Restricting this exponential map $S\times[-\frac{1}{2},\frac{1}{2}]$ gives a codimension 0 submanifold $Q_{p}^{0}\subseteq Q_{p}$ which is $2$-bilipschitz to $S\times[-\frac{1}{2},\frac{1}{2}]$ and separates the rest of $Q_{p}$ into two components $Q_{p}^{+}$ (above) and $Q_{p}^{-}$ (below).
Inside $S$, we choose a net of $\sim 2^{2p}$ balls $\{B_{i}\}$ of radius $2^{-\frac{p}{n-1}}$. We then define a map to $C$ which sends $S\times[-\frac{1}{2},\frac{1}{2}]$ to $N_{p}$:
$f(x,t)=\bigl{(}\rho\bigl{(}p+t+{\textstyle\frac{1}{2}}\bigr{)}\theta(x),p+t+{\textstyle\frac{1}{2}}\bigr{)},$
where $\theta(x)=(1,0,\ldots,0)$ for $x$ outside the $B_{i}$ and $\theta$ maps each $B_{i}$ homeomorphically to $S^{n-1}\setminus(1,0,\ldots,0)$. This map can now be extended to $Q_{p}$ so that $Q_{p}^{-}$ is mapped to the neck at $p$ and $Q_{p}^{+}$ is mapped to the neck at $p+1$, with degree $\sim 2^{2p}$ on the upper and lower boundary. Note that this relies on the fact that the cross-section of the neck has area $2^{-p}$ times the cross-sectional area of the center of a nodule; this allows all the homeomorphic preimages of the neck to be routed through the space outside $Q_{p}^{0}$.
One needs to do some further routing to match the map in the middle section of the bottom surface of $Q_{p}$ with that on the top section of $Q_{p-1}$, but again it’s not hard to arrange for enough space to do this, perhaps at the expense of a constant increase in the Lipschitz constant.
See Figure 2 for an illustration.
[…]
Figure 2. A schematic illustration of $Q_{p}$, with the embedding of $S\times[-\frac{1}{2},\frac{1}{2}]$ and several preimages of the corresponding nodule shown.
5.3. Proof of Theorem 1.9
Now we prove that elliptic manifolds satisfying certain hypotheses must have nilpotent fundamental group. Effectively, these are manifolds over which any wrapping map equidistributes, that is, it must hit every point roughly equally often. This condition holds for quasiregular maps to any manifold, so it is not surprising that it shows up here.
Theorem 1.9.
Suppose that $M$ is a Riemannian $n$-manifold of finite volume, and suppose that the volume form of $M$ can be written as $\omega+d\alpha$ where $\omega$ is compactly supported and $\alpha$ is bounded. If $M$ is elliptic, then $\pi_{1}(M)$ is nilpotent and any finitely generated subgroup of it has growth $O(R^{n})$.
Note that $\pi_{1}(M)$ may be an infinitely generated group, for example $\mathbb{Q}$. The hypothesis holds for example for manifolds of pinched negative curvature as well as real analytic manifolds of nonpositive curvature, including locally symmetric spaces. The proof in both cases is the same and follows from results described in [4, §10, 13].
If $M$ is a Riemannian $n$-manifold of finite volume (not necessarily complete), the Cheeger isoperimetric constant of $M$ is defined by
$h(M)=\inf\left\{\frac{\operatorname{vol}(\partial A)}{\min\{\operatorname{vol}A,\operatorname{vol}A^{c}\}}:A\subset M\text{open}\right\}.$
When $M$ is noncompact, an equivalent (and more commonly used) definition is
$h(M)=\inf\left\{\frac{\operatorname{vol}(\partial A)}{\operatorname{vol}A}:A\text{ is a relatively compact open subset with }\operatorname{vol}A\leq\frac{1}{2}\operatorname{vol}M\right\}.$
To see the equivalence, note that in the first definition one can make sure that the smaller subset is relatively compact (without changing the volumes too much) by intersecting it with a large ball around some base point; the coarea formula ensures that one can pick the radius of the ball so that $\operatorname{vol}(\partial A)$ is affected an arbitrarily small amount.
The Cheeger constant of a manifold can be zero, for example for those constructed in Section 5.2. In future work, we will show that $M$ has nonzero Cheeger constant if and only if it satisfies the hypothesis in Theorem 1.9. This equivalence has an analogue for manifolds of infinite volume stated by Gromov [12, §4.1] which can be proved using similar techniques.
Proof of Theorem 1.9..
Let $f:\mathbb{R}^{n}\to M$ be a wrapping map.
By Theorem LABEL:IP-dual, we can write $d\operatorname{vol}_{M}=\omega+d\alpha$, where $\omega$ is compactly supported and $\alpha$ is bounded. Then by Stokes’ theorem,
$\int_{B(R)}f^{*}\omega=\int_{B(R)}f^{*}d\operatorname{vol}_{M}-\int_{\partial B(R)}f^{*}\alpha,$
and therefore
$\left\lvert\int_{B(R)}f^{*}\omega\right\rvert\geq\left\lvert\int_{B(R)}f^{*}d\operatorname{vol}_{M}\right\rvert-O(R^{n-1}).$
In particular, suppose $\omega$ is supported on a compact set $K$. Then
$\limsup_{R\to\infty}\operatorname{vol}(B(R)\cap f^{-1}(K))\geq\varepsilon R^{n}$
for some $\varepsilon>0$.
Now let $K^{\prime}\supset K$ be another compact set which contains the $r$-neighborhood of $K$; we can assume $K^{\prime}$ is a smooth codimension zero submanifold of $M$ and give it a triangulation $\mathcal{T}$. Let $\lambda>0$ be such that every $\lambda$-ball in $K^{\prime}$ is contained in the star of some vertex of $\mathcal{T}$. Choose a triangulation $\mathcal{T}^{\prime}$ of $B(R)$ whose simplices are $C(n)$-bilipschitz to a linear simplex of side length $\min\{r,\lambda\}$ and have diameter $<\min\{r/2,\lambda/2\}$. Then there is a subcomplex $X\subseteq\mathcal{T}^{\prime}$ which contains the image of $K$ and whose star lies in $K^{\prime}$. Moreover, by the usual proof of the simplicial approximation theorem, see e.g. [14], a perturbation of $f$ gives a map $f^{\prime}$ whose restriction to $X$ is simplicial on $\mathcal{T}$.
We can assume that the induced map $\pi_{1}(K)\to\pi_{1}(M)$ is injective. Now consider the induced map from the dual graph of $X$ to the dual graph $\Gamma$ of $\widetilde{T}$, the universal cover of $T$. This defines a 00-cochain in $\Gamma$ of volume $\sim R^{n}$ whose coboundary has volume $\sim R^{n-1}$. Since $\Gamma$ is quasi-isometric to the Cayley graph of $\pi_{1}(K)$, it has the same isoperimetric and growth functions. Therefore, by the Varopoulos isoperimetric inequality (see e.g. [13, §6.E]), $\pi_{1}(K)$ has growth rate $O(R^{n})$.
Now we could have taken $\pi_{1}(K)$ to be any finitely generated subgroup of $\pi_{1}(M)$. Therefore the nilpotency class of any such subgroup is at most $n-1$ (in fact, at most around $\sqrt{n}$, but this is harder to see), and so is the nilpotency class of $\pi_{1}(M)$. ∎
5.4. Proof of Theorem 5.1
Lemma 4.7 shows that the theorem holds for $n=2$. We will use the intermediate map $f_{0}$ defined in that lemma to construct maps $\mathbb{R}^{n}\to S^{n}\setminus S^{n-2}$ of positive asymptotic degree for $n\geq 3$.
First define a Lipschitz map of positive asymptotic degree $g_{n-2}:\mathbb{R}^{n-2}\to S^{n-2}$; for example, we can compose the universal covering map $\mathbb{R}^{n-2}\to T^{n-2}$ with a map that collapses the $(n-3)$-skeleton of the torus.
We can view $S^{n}$ as a quotient
$S^{n}\cong S^{1}*S^{n-2}\cong(D^{2}\times S^{n-2})/((x,y)\sim(x,y^{\prime})\text{ for }x\in\partial D^{2},\,y,y^{\prime}\in S^{n-2}).$
(Here * represents the join of two spaces.) Denote this quotient map by $q:S^{n-2}\times D^{2}\to S^{n}$. Finally let $\rho:\mathbb{R}^{n-2}\to\mathbb{R}^{n-2}$ be an orientation-reversing map (say, flipping one coordinate).
Now we define our desired map similarly to the way $f$ was defined above:
$f_{n}(x,y,\mathbf{z})$ $=q(f_{0}(x,y),g_{n-2}(\mathbf{z}))$ $(x,y)$ $\in[-1,1]\times[0,\infty)\times\mathbb{R}^{n-2}$
$f_{n}(x,y,\mathbf{z})$ $=q(f_{0}(2-x,y),g_{n-2}\rho(\mathbf{z}))$ $(x,y)$ $\in[1,3]\times[0,\infty)\times\mathbb{R}^{n-2}$
$f_{n}(x+4,y,\mathbf{z})$ $=f_{n}(x,y,\mathbf{z})$
$f(x,-y,\mathbf{z})$ $=f_{n}(x,y,\rho(\mathbf{z})).$
This map is continuous since on the boundaries between subdomains, the map depends only on $x$ and $y$. Again, a computation shows that $f_{n}$ is Lipschitz, has the same orientation at every regular point, and that its Jacobian is bounded away from zero on a large subset. This implies that it is of positive asymptotic degree.
Acknowledgements
The authors would like to thank Ben Lowe for comments drawing attention to the parallels between their previous work, which sparked this collaboration. F.M. would also like to thank Pekka Pankka and Susanna Heikkilä for useful conversations and comments.